### v1

# Why Quantum-like Models of Cognition Work

Abstract
It is shown that Brownian motions executed by state points of neural membranes generate a Schrödinger-like equation with $\hbar/m$ replaced by the coefficient of diffusion $\sigma$ of the substrates.
One of the fundamental problems of the entire body of pioneering work on quantum-like models of the social sciences (see Refs [1, 2, 3] and references therein) is that in spite of its successes, it is plagued by the absence of any empirical evidence of the quantum-like nature of the brain as well as by the interpretational problems associated with quantum mechanics. The brain is an open macroscopic hot and and noisy system in which decoherence should occur almost instantaneously. Nevertheless, there is strong evidence that classical probability theory fails to account for a number of observed psychological behaviour like the order effect [4, 5], the conjunction effect [6] and the disjunction effect [7, 8] which are all indicative of quantum-like behaviour.
The purpose of this paper is to show that it is precisely the stochastic nature of the neural processes underlying cognition that generates a Schrödinger-like equation with the factor $\hbar/m$ replaced by the square root $\sigma$ of the diffusion coefficient of the relevant neural substrates. This is in the spirit of Nelson [9, 10, 11] who showed that quantum mechanics is an emergent theory, resulting from underlying stochastic processes (Brownian motion) that are entirely classical. Similarly, the quantum-like behaviour of cognition can be traced to classical Brownian motion in neural substrates.
There is a whole body of empirical and theoretical work which has established the stochastic behaviour of such substrates [12]. In particular, random walk and diffusion models for spike activity of a neuron have been proposed [13, 14]. Stated in the language of neurophysiology, the electrical state of polarization of the somatic and dendritic membrane may be represented by a “state point” which executes random walk. Each incoming elemental EPSP (Excitatory Post Synaptic Potential) moves the state point one unit toward the threshold of neuron firing, and each incoming elemental IPSP (Inhibitory Post Synaptic Potential) moves the state point one unit away from the threshold. If the average rate of incoming elemental EPSP and elemental IPSP are the same, there is an equal probability at any time that the state point moves either a unit toward or a unit away from the threshold, i.e. there is no “bias toward” either input. Immediately after the state point has attained the threshold and caused the production of an action potential, it returns to the resting potential, only to begin its random walk again. Such a model would be a simple random walk model.
In reality, however, these two rates may sometimes be different, and in realistic physiological models it would be far more reasonable to assume that that there is some excess of either EPSP or IPSP input. In this case, the probability for the state point to move one unit toward the threshold will be different from the probability for it to move away from the threshold. Considered as a diffusion process, the difference between these probabilities can be considered a“drift velocity”, either toward or away from the threshold. It is this “random walk with drift” [13] which is of particular interest to quantum-like cognitive modelling.
Following Nelson, let us consider the following model.
(A) Let the “state point” of a neuron be specified by a single number $X(t)$, the membrane potential. As time passes and the electrical state of the membrane varies, the state point will move back and forth along a straight line, executing Brownian motion without friction. This means the ‘motion’ is a conservative diffusion process, i.e. a process in which there is on the average no exchange of energy between the state point and its neural background, i.e. the mutual exchanges of energy average out to zero but are responsible for “quantum-like fluctuations” of the system. The motion is described by a Markoff process in the state space.
To describe the spontaneous spiking of neurons, one must introduce a resting potential and an absorbing barrier, the threshold. If at any time the state point reaches the threshold, the neuron produces an action potential.
(B) The path of a state point driven by Brownian motion is written as the Itô stochastic differential equation (SDE)
$dX(t)=b_{f}(X(t),t)dt+\sigma dW_{f}(t)$ (1)
where $b_{f}$ is the forward drift velocity (caused by EPSP) which depends on the current position $x=X(t)$, $\sigma$ is the square root of the diffusion coefficient and $dW_{f}(t)$ is a forward Wiener process. Since these processes are conservative, backward processes (caused by IPSP) exist,
$dX(t)=b_{b}(X(t),t)dt+\sigma dW_{b}(t)$ (2)
where $dW_{b}(t)$ is the backward Wiener process.
(C) The diffusion coefficient $\sigma^{2}$ is determined by the physiological characteristics of the membrane.
(D) The solutions $X(t)$ of the stochastic equations are known to be continuous at all state points but nowhere differentiable. Hence, Nelson suggested the following average forward and backward differentials which we adopt:
$D_{f}X(t)$ $=$ $\lim_{\Delta t\rightarrow 0}E_{t}\left[\frac{X(t+\Delta t)-X(t)}{\Delta t}\right],$ (3)
$D_{b}X(t)$ $=$ $\lim_{\Delta t\rightarrow 0}E_{t}\left[\frac{X(t)-X(t-\Delta t)}{\Delta t}\right]$ (4)
where $E_{t}$ denotes the expectation conditional on $X(t)=x$. For differentiable curves $D_{f}X(t)=D_{b}X(t)=\dot{x}=v(t)$, the ‘velocity’ of the state point. Let us postulate Newton’s law for the stochastic acceleration,
$ma(X(t))=m\frac{1}{2}(D_{f}D_{b}+D_{b}D_{f})X(t)=F(X(t))$ (5)
where $m$ denotes state inertia, the ability of biological systems to keep a functional state at rest or in activity and is an active process of resistance to change in state. It follows that the drift coefficients in the forward and backward equations are given by
$D_{f}(X(t))=b_{f}(X(t),t),\,\,\,D_{b}(X(t))=b_{b}(X(t),t)$ (6)
This amounts to a complete description of the motion, as in classical mechanics.
The forward and backward stochastic differential equations are equivalent to two Fokker-Planck equations
$\frac{\partial}{\partial t}\rho(x,t)$ $=$ $-\frac{\partial}{\partial x}\left[b_{f}(x,t)\rho(x,t)\right]+\frac{\sigma^{2}}{2}\frac{\partial^{2}}{\partial x^{2}}\rho(x,t),$ (7)
$\frac{\partial}{\partial t}\rho(x,t)$ $=$ $-\frac{\partial}{\partial x}\left[b_{b}(x,t)\rho(x,t)\right]-\frac{\sigma^{2}}{2}\frac{\partial^{2}}{\partial x^{2}}\rho(x,t)$ (8)
where $\rho(x,t)$ is the probability density of the random variable $X(t)$. Addition of these two equations results in the continuity equation
$\frac{\partial}{\partial t}\rho(x,t)+\frac{\partial}{\partial x}[v(x,t)\rho(x,t)]=0$ (9)
with the current velocity $v(x,t)=(b_{f}(x,t)+b_{b}(x,t))/2$. The difference of the forward and backward drifts $u(x,t)=(b_{f}(x,t)-b_{b}(x,t))/2$ is the osmotic velocity. Subtracting the two Fokker-Planck equations we get
$u(x,t)=\frac{\sigma}{2}\frac{\partial}{\partial x}\ln[\rho(x,t)]=\frac{\sigma}{2}\frac{\partial_{x}\rho}{\rho}=\sigma\frac{\partial R}{\partial x}$ (10)
where $\ln\rho(x,t)=2R(x,t)$. The coupled forward-backward stochastic differential equations for the position process can thus be written as
$dX(t)$ $=$ $\left(v(X(t),t)+u(X(t),t)\right)+\sigma dW_{f}(t),$ (11)
$dX(t)$ $=$ $\left(v(X(t),t)-u(X(t),t)\right)+\sigma dW_{b}(t).$ (12)
It follows from this that the current velocity is curl-free and can be written as
$v(x,t)=\frac{1}{m}\frac{\partial}{\partial x}S(x,t)$ (13)
where $S(x,t)$ is a scalar function which can be identified with the action.
Now, following Guerra and Morato [15], let us introduce the Lagrangian field
${{\cal{L}}}=\frac{1}{2}m(v^{2}-u^{2})(x,t)-V(x)$ (14)
where $V(x)$ is the electrostatic potential, from which the action $S(x,t)$ can be constructed. It can then be shown, using the variational principle, that the main features of Nelson’s stochastic mechanics including eqns (10) and (13) can be derived from such an action.
Using stochastic control theory, Guerra and Morato also derived the following differential equations for the functions $R$ and $S$:
$\frac{\partial S}{\partial t}+\frac{1}{2m}\left(\frac{\partial S}{\partial x}\right)^{2}+V+V_{Q}$ $=$ $0,\,\,\,V_{Q}=-\frac{m\sigma^{2}}{2}\left[\left(\frac{\partial R}{\partial x}\right)^{2}+\frac{\partial^{2}R}{\partial x^{2}}\right],$ (15)
$\frac{\partial R}{\partial t}+\frac{1}{2m}\left(R\frac{\partial^{2}S}{\partial x^{2}}+2\frac{\partial R}{\partial x}\frac{\partial S}{\partial x}\right)$ $=$ $0$ (16)
The first equation is the Hamilton-Jacobi-Bellman equation, i.e. the Hamilton-Jacobi equation with an additional stochastic term $V_{Q}$ which takes the form
$V_{Q}=-\frac{m\sigma^{2}}{4}\left[\frac{\partial_{x}^{2}\rho}{\rho}-\frac{(\partial_{x}\rho)^{2}}{2\rho^{2}}\right]$ (17)
in terms of $\rho=e^{2R}$. It is the analog of the Bohm quantum potential. The second equation can also be written in terms of $\rho$ as
$\frac{\partial\rho}{\partial t}+\partial_{x}\left[\rho\frac{\partial_{x}S}{m}\right]=0$ (18)
which, using eqn (13) for the current velocity, is a continuity equation. These two coupled partial differential equations determine the stochastic process. These equations can be derived from the Schrödinger-like equation
$im\sigma\frac{\partial}{\partial t}\psi(x,t)=\left(-\frac{m\sigma^{2}}{2}\partial_{x}^{2}+V(x)\right)\psi(x,t)$ (19)
by putting $\psi=\exp(R+iS)=\sqrt{\rho}\exp(iS/m\sigma)$ and separating the real and imaginary parts [15, 16]. The coefficient $\sigma$ clearly plays the role of the factor $\hbar/m$ in quantum mechanics.
The wave function $\psi$ describes the Markov process completely:
$\rho$ $=$ $|\psi|^{2},$ (20)
$u$ $=$ $\sigma\partial_{x}\Re\ln\psi,$ (21)
$v$ $=$ $\sigma\partial_{x}\Im\ln\psi.$ (22)
This is the ‘Nelson map’. It maps the probability distribution function and the current and osmotic velocities in the neural substrates to the wave function. In other words, it associates a diffusion process in the substrates to every solution of the Schröodinger-like equation (19).
It should be mentioned that entangled states exist in stochastic mechanics and that stochastic mechanics and quantum mechanics agree in predicting all the observed correlations at different times. The reader is referred to the papers by Faris [17] and Petroni and Morato [18] for details. It should also be pointed out that the violation of Bell’s locality is an intrinsically statistical phenomenon (it cannot produce superluminal signals!) and hence a stochastic framework should be the most appropriate for facing such a challenging problem.
Finally, it is well known that stochastic mechanics and Feynman’s path integral formulation of quantum mechanics are intimately related [19, 20]. Feynman’s path integral has been applied to finance [21] with some success. Hence a stochastic mechanics approach to quantum finance should also be possible.